Kay spends 250 minutes per week exercising. Her ratio of aerobics to weight training is 7 to 5. How many minutes per week does she spend on aerobics? How many minutes per week does she spend on weight training?
Answer by stanbon(75887)

Kay spends 250 minutes per week exercising. Her ratio of aerobics to weight training is 7 to 5. How many minutes per week does she spend on aerobics? How many minutes per week does she spend on weight training?
------------
aero/weight = 7/5 = 7x/5x
--------
EQUATION:
7x+5x=250
12x=250
x=20.833333
aerobics time = 7x = 7(20.8333333...) = 145.83333... minutes
Cheers,
Stan H.